OpenStudy (anonymous):

5x+4=-11 What can I do farther to isolat the variable? Cann I add 4 or Subtract 4 from both sides or can I Subtract 5x from bth sides? Which one (:

OpenStudy (aravindg):

subtract 4 and div by 5

OpenStudy (anonymous):

subtract 4 from both sides

OpenStudy (anonymous):

1.Add 4 2.Subtract 4 3.Subtract 5x

OpenStudy (anonymous):

2 (:

OpenStudy (anonymous):

5x+4=-11 you can subtract four from both sides 5x+4=-11 -4 -4 You would get this: 5x = -14 Then if you want to get x=? Then you do this 5x/5 = -14/5 x = -2.8